OpenStudy (anonymous):

6(x-1)=36 please help. include steps.

OpenStudy (bahrom7893):

6(x-1) = 36 6*(x-1) = 6*6 (x-1) = 6 x-1 = 6 x-1 = 7-1 x = 7

OpenStudy (bahrom7893):

Please tell me that makes sense

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lolkimkim

OpenStudy (swissgirl):

6(x-1) = 36 \( \large {6(x-1) \over 6} = { 36 \over 6}\) (x-1) = 6 x-1+1 = 6+1 x= 7